Firms HL and LL are identical except for their financial leverage ratios and the interest rates they pay on debt. Each has $10 million in invested capital, has $2 million of EBIT, and is in the 40% federal-plus-state tax bracket. Firm HL, however, has a debt-to-capital ratio of 60% and pays 11% interest on its debt, whereas LL has a 20% debt-to-capital ratio and pays only 10% interest on its debt. Neither firm uses preferred stock in its capital structure.

  1. Calculate the return on invested capital (ROIC) for each firm. Round your answers to two decimal places.

    ROIC for firm LL: _____ %
    ROIC for firm HL:  _____ %
  2. Calculate the rate of return on equity (ROE) for each firm. Round your answers to two decimal places.

    ROE for firm LL:   _____ %
    ROE for firm HL:   _____ %
  3. Observing that HL has a higher ROE, LL's treasurer is thinking of raising the debt-to-capital ratio from 20% to 60% even though that would increase LL's interest rate on all debt to 15%. Calculate the new ROE for LL. Round your answer to two decimal places.

    _____ %
Answer #1

a.ROIC = EBIT(1-T)/Capital

LL = 2 million(1-40%)/10 million = 12%

HL = 2 million(1-40%)/10 million = 12%

b.Return on Equity = Income for Equity/Equity Funds

LL = (2 million – 2*10%)(1-40%)/8

= 13.5%

HL = (2 – 0.66)(1-40%)/4

= 20.1%

c.ROE = (2 – 0.9)(1-40%)/4

= 16.5%